the sum of 4 consecutive integers of 114
Answer by jim_thompson5910(35256)

n + n+1 + n+2 + n+3 = 114

4n+6=114

4n=114-6

4n=108

n=(108)/(4)

n=27


So the first number is 27 and the four numbers are: 27, 28, 29, 30